Suppose you want to have $300,000 for retirement in 30 years. Your account earns 10% interest. a) How much would you need to deposit in the account each month? $ b) How much interest will you earn?
Answer #1

(1+5.) - 1 FV = PMT

FV=300000

t=30 years

r=10% = 0.1

n=12 for monthly deposite

2 0 1 300000 = PMT

300000 = PMT (1 +0.008333,360 0.008333

300000 = PMT (1.008333300 - 1 0.008333

300000 = PMT 19.83503 - 1 0.008333

300000 = PMT 18.83503 0.008333

300000 = PMT 2260.48792

PMT = 300000 [2260.48792]

PMT = 132.71................monthly deposit

.

.

.

.

interest is given by

1=FV - PMT.n.t

I=300000-132.71471\cdot 12\cdot 30

I = 300000 - 47777.2956

I=252222.7044

I 252222.7................interest